# RobustFormer: Noise-Robust Pre-training for Images and Videos

### v1

@@ 4.2 Training
For video datasets, we apply the same foundational framework with adaptations tailored to temporal data. Pretraining on UCF-101 and HMDB-51 involves 800 epochs using 16-frame sequences, a batch size of 8, a tube masking ratio of 0.9, a sampling rate of 4, and a decoder depth of 4. The same node configuration is used as in the image tasks. Fine-tuning for these datasets is conducted over 100 epochs. For larger-scale datasets like Kinetics-400P, training is distributed across 12 nodes with V100 GPUs (2 GPUs and 40 cores per node) for 400 epochs, maintaining a tube masking ratio of 0.9, sampling rate of 4, and decoder depth of 4 but with an increased batch size of 16. We use DeepSpeed during video fine-tuning to manage larger batch sizes efficiently.
Our evaluation focuses on the model’s robustness to real-world noise, including shot noise, rain noise, Gaussian noise, packet loss, speckle noise, and tampering noise, as identified in prior benchmarks [37, 57]. Unlike traditional methods, we do not introduce noise augmentations during training or validation, enabling us to assess the model’s ability to generalize to unpredictable noise patterns at varying intensities. This approach mirrors real-world conditions, where noise is often diverse and unstructured. For instance, practical noise types like rain, critical for autonomous driving, and packet loss, which may affecting video-based applications. Additionally, testing on image datasets shows that our DWT-based approach is not only robust but also achieves a 2.5% accuracy improvement on clean datasets.
@@ 4.3 Evaluation
To measure video robustness we use two metrics for relative and absolute accuracy drops. After training model , we first compute the accuracy on the clean set and accuracy for perturbation and severity . The absolute robustness and relative robustness are then computed as and . The aggregated performance for all models can be thus achieved by averaging across severity levels to get and . Table 4 shows and , which are averaged and across the perturbation categories for Kinetics-400 dataset. Our RF variants even without augmented training demonstrated significant robustness compared to some of the methods that used augmentation during training. Similarly, Table 5 shows the relative robustness scores for individual perturbations averaged across 5 severity levels. RF models here demonstrated better relative robustness to recent state-of-the-art methods. Moreover, top-1 and top-5 accuracy performance on additional noise types like rain, packetloss for UCF-101 dataset are shown in Figure 5, where RF methods are significantly superior especially in noise with high severity levels.
@@ 4.4 Ablation Study
Our preferred RobustFormer variant, RF-AA, demonstrates competitive and, in many cases, superior performance in both top-1 and top-5 accuracy compared to other configurations. Variants RF-I and RF-IA represent RobustFormer with the inclusion of the IDWT step and with both IDWT and attention steps, respectively. The IDWT step, commonly employed in traditional DWT-based methods, complements the regular DWT averaging technique.
Through our experiments, we show that incorporating the proposed DWT attention mechanism significantly enhances performance, particularly for complex noise types. For example, while the average filter alone struggles with salt-and-pepper noise, the addition of the DWT attention module leads to a marked improvement in accuracy. Similarly, the DWT attention module boosts accuracy in certain noise conditions even when applied to IDWT-based configurations, as highlighted in Figure 6 (b-d). This demonstrates the critical role of our dwt based attention mechanism in improving robustness against challenging noise scenarios.
@@ 5 Discussion and Conclusion
Noise-robust models are essential for many real-world applications, where unpredictable and unique noise patterns can significantly impact performance. While some deep learning models rely on heavy augmentation and others on rule-based filtering, these approaches are limited in scalability due to the inherent randomness and diversity of noise. RobustFormer addresses this challenge by utilizing transformers’ adaptability and the noise-handling capabilities of Wavelet Transform. This approach enables efficient decomposition of spatio-temporal information, selectively focusing on low-frequency components to enhance robustness across diverse noise types without the need for exhaustive augmentation or complex filtering rules thus making few assumptions about the evaluation noises. We validated our method on both image and video benchmark datasets with benchmark noise types for both and show our model’s significance compared to prior works.

### v2

@@ 4.2 Training
For video datasets, we apply the same foundational framework with adaptations tailored to temporal data. Pretraining on UCF-101 and HMDB-51 involves 800 epochs using 16-frame sequences, a batch size of 8, a tube masking ratio of 0.9, a sampling rate of 4, and a decoder depth of 4. The same node configuration is used as in the image tasks. Fine-tuning for these datasets is conducted over 100 epochs. For larger-scale datasets like Kinetics-400P, training is distributed across 12 nodes with V100 GPUs (2 GPUs and 40 cores per node) for 400 epochs, maintaining a tube masking ratio of 0.9, sampling rate of 4, and decoder depth of 4 but with an increased batch size of 16. We use DeepSpeed [40] during video fine-tuning for memory-saving optimizations.
Our evaluation focuses on the model’s robustness to real-world noise, including shot noise, rain noise, Gaussian noise, packet loss, speckle noise, and tampering noise, as identified in prior benchmarks [37, 57]. These pertubations mirror real-world conditions, where noise is often diverse and unstructured. For instance, practical noise types like rain, critical for autonomous driving, and packet loss, which may affecting video-based applications. Additionally, testing on image datasets shows that our DWT-based approach is not only robust but also achieves a 2.5% accuracy improvement on clean datasets.
@@ 4.3 Evaluation
To measure video robustness we use two metrics for relative and absolute accuracy drops. After training model , we first compute the accuracy on the clean set and accuracy for perturbation and severity . The absolute robustness and relative robustness are then computed as and . The aggregated performance for all models can be thus achieved by averaging across severity levels to get and . Table 4 shows and , which are averaged and across the perturbation categories for Kinetics-400 dataset. Our RF variants even without augmented training demonstrated significant robustness compared to some of the methods that used augmentation during training. Similarly, Table 6 shows the relative robustness scores for individual perturbations averaged across 5 severity levels. RF models here demonstrated better relative robustness to recent state-of-the-art methods. Moreover, top-1 and top-5 accuracy performance on additional noise types like rain, packetloss for UCF-101 dataset are shown in Figure 5, where RF methods are significantly superior especially in noise with high severity levels. Additional results (including for ImageNet-Tiny-200 dataset evaluated on our custom noise) are kept in supplementary.
@@ 4.4 Ablation Study
We ablate the effects of the IDWT step and the proposed DWT-based attention module across different RobustFormer variants. Our preferred model, RF-AA, achieves competitive and often superior top-1 and top-5 accuracy compared to all other configurations. The variants RF-I and RF-IA correspond to adding the IDWT reconstruction step and adding both IDWT and DWT-attention, respectively. While the IDWT step is commonly used in traditional DWT pipelines, it offers only limited gains when used alone.
Our experiments show that the DWT-attention module is the primary source of robustness improvements. Averaging alone struggles with high-impulse corruptions such as salt-and-pepper noise, whereas adding DWT attention yields substantial accuracy gains. Moreover, DWT attention continues to improve performance even within IDWT-based variants (Fig. 6b–d). Removing the IDWT step also reduces computational cost by 7.1 GFLOPs when comparing RF-OA with RF-IA (Table 6). Overall, these ablations highlight the central role of DWT-based attention in achieving strong robustness under diverse noise conditions.
@@ 5 Discussion and Conclusion
Noise-robust models are essential for many real-world applications, where noise patterns can significantly impact performance. While some deep learning models rely on heavy augmentation and others on rule-based filtering, these approaches are limited in scalability due to the inherent randomness and diversity of noise. Our approach enables efficient decomposition of spatio-temporal information, selectively focusing on low-frequency components to enhance robustness across diverse noise types without the need for exhaustive augmentation or complex filtering rules, thus making few assumptions about the evaluation noises. We validated our method on both image and video benchmark datasets with benchmark noise types for both and show our model’s significance compared to prior works.